OpenStudy (anonymous):

Look at the table of values below. x y 1 -1 2 -3 3 -5 4 -7 Which equation is represented by the table? Select one of the options below as your answer: A. y = 1 - 2x B. y = -x- 1 C. y = x- 2 D. y = 2x- 1

OpenStudy (raden):

check for the first one y = 1 - 2x if x = 1 then y = 1 - 2(1) = 1 - 2 = -1 (true) if x = 2 then y = 1 - 2(2) = 1 - 4 = -3 (true) if x = 3 then y = 1 - 2(3) = 1 - 6 = -5 (true) ... ... so on

OpenStudy (anonymous):

so whats the answer?

OpenStudy (raden):

i think u can take a conclusion based information above

OpenStudy (anonymous):

im just a 9th grader trying to pass , math is hard for me i pay attention but i dont get it...

OpenStudy (anonymous):

so the answer is D?

OpenStudy (raden):

y = 1 - 2x